# A fast Bellman algorithm for the real Monge–Ampère equation

### v2

@@ 5. Description of our algorithm
Next, we construct a sequence of functions $u_{k}$ where $u_{k}$ solves the discretized version of the following Dirichlet problem, using a narrow stencil second order finite difference solver:
@@
We terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or when the number of iterations exceeds 10 0001000010\,00010 000.
@@
which is then solved using Gauss–Seidel iteration, keeping the boundary values fixed and updating one $u_{ij}$ at the time, using (5.4). We should note that this is computationally expensive, especially in our NumPy implementation where explicit loops are slow, so M1, and to some extent our proposed method, are the ones that would benefit the most from an optimized implementation.
In our comparison we chose the same termination criterion for all methods, i.e. to terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or the number of iterations exceeds 10 0001000010\,00010 000 (for M1, we allow up to 300 000300000300\,000300 000 iterations, where one iteration means updating each $u_{ij}$ once).
@@ 6.1.2. Degenerate example with regularisation
where ϵ+italic-$\epsilon\in\mathbb{R}_{+}$, and the right-hand side of the equation is
@@ 6.2.2. Degenerate example
$N$ iter time (s) iter time (s) iter time (s)
@@ 6.2.3. Constant Monge–Ampère
Table 2. Minimum value for the computed solution to the constant Monge–Ampère example. Method M1 and M2 produce the same values (to four points of accuracy). The computations with the M2 method were terminated after 10 0001000010\,00010 000 iterations, and the computations with M1 was terminated after 300 000300000300\,000300 000 iterations, marked by * in the table. The results for the wide stencil finite difference discretizations (4th–6th column) are taken from Oberman [26].
$N$ Bellman M1, M2 9pt stencil 17pt stencil 33pt stencil
@@
Table 3. Running time for the constant Monge–Ampère example. Computations marked by * were terminated after 10 0001000010\,00010 000 (for method M2) and 300 000300000300\,000300 000 (for M1) iterations.
@@
$N$ iter time (s) iter time (s) iter time (s)
@@
$41$ $19$ $0.5$ $5086$ $14.3$ $1539$ $23.2$
$61$ $19$ $1.6$ 11 1461114611\,14611 146 $69.9$ $3224$ $150.3$
$81$ $21$ $3.8$ 19 3791937919\,37919 379 $222.5$ $5443$ $546.8$
$101$ $19$ $5.9$ 29 6982969829\,69829 698 $518.3$ $8158$ $1513.7$
$127$ $20$ $10.8$ 46 1524615246\,15246 152 $1297.7$
$255$ $29$ $102.7$ 174 401174401174\,401174 401 20 896.020896.020\,896.020 896.0 * *
$511$ $25$ $870.7$
Running the Bellman algorithm on Equation (6.2) produces a function that (barely) fails to be convex near the corners of $\Omega$, but the same is true for Method M2, see Figure 6. This example was also studied by Dean and Glowinski [10] (using a variational formulation) and Oberman [26] (wide stencil finite difference discretization), all with similar results. In [5], they suggest comparing the different methods using the minimum (i.e. the value at the origin) of the solution. The wide stencils methods from Oberman [26] are known to converge monotonically to a viscocity supersolution of Equation (2.1), and are thus too large. If the solution produced by the Bellman method had been convex, Proposition 3.3 would have implied that it too is too large, but since convexity fails near the corners of the square, we have no guarantee for this. Our method produces a solution which is slightly larger than the M2 and M1 methods, see Table 2.
@@ 6.3.2. Unbounded Monge–Ampère
If we instead do the computations on a slightly smaller square, e.g. $[0,0.99]^{2}$, the three methods run at full accuracy ($\sim N^{-2}$), and the Bellman method runs 2–3 times faster than M2.
@@ 7. Conclusions
[…]

### v3

@@ 5. Description of our algorithm
Next, we construct a sequence of functions $u_{k}$ where $u_{k}$ solves the discretized version of the following Dirichlet problem, using a central difference second order finite difference solver:
@@
We terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or when the number of iterations exceeds 10 00010\,000.
@@
which is then solved using Gauss–Seidel iteration, keeping the boundary values fixed and updating one $u_{ij}$ at the time, using (5.4). We should note that this is computationally expensive, especially in our NumPy implementation where explicit loops are slow, so M1, and to some extent our proposed method, are the ones that would benefit the most from an optimized implementation. Here, it would be natural to update the whole matrix of functions values at once, using (5.4) instead of doing it element by element, but apparently this destroys the convergence to the actual solution of the Monge–Ampère equation.
In our comparison we chose the same termination criterion for all methods, i.e. to terminate the computation when $\|u_{k}-u_{k-1}\|_{\infty}<10^{-12}$ or the number of iterations exceeds 10 00010\,000 (for M1, we allow up to 300 000300\,000 iterations, where one iteration means updating each $u_{ij}$ once).
@@ 6.1.2. Degenerate example with regularisation
where $\epsilon\in\mathbb{R}_{+}$, and the right-hand side of the equation is
@@ 6.2.2. Degenerate example
NN iter time (s) iter time (s) iter time (s)
@@ 6.2.3. Constant Monge–Ampère
Table 2. Minimum value for the computed solution to the constant Monge–Ampère example. Method M1 and M2 produce the same values (to four points of accuracy). The computations with the M2 method were terminated after 10 00010\,000 iterations, and the computations with M1 was terminated after 300 000300\,000 iterations, marked by * in the table. The results for the wide stencil finite difference discretizations (4th–6th column) are taken from Oberman [27].
NN Bellman M1, M2 9pt stencil 17pt stencil 33pt stencil
@@
Table 3. Running time for the constant Monge–Ampère example. Computations marked by * were terminated after 10 00010\,000 (for method M2) and 300 000300\,000 (for M1) iterations.
@@
NN iter time (s) iter time (s) iter time (s)
@@
$41$ 1919 $0.5$ $5086$ $14.3$ $1539$ $23.2$
$61$ 1919 $1.6$ 11 14611\,146 $69.9$ $3224$ $150.3$
$81$ 2121 $3.8$ 19 37919\,379 $222.5$ $5443$ $546.8$
$101$ 1919 $5.9$ 29 69829\,698 $518.3$ $8158$ $1513.7$
$127$ 2020 $10.8$ 46 15246\,152 $1297.7$ * *
$255$ $29$ $102.7$ 174 401174\,401 20 896.020\,896.0 * *
$511$ $25$ $870.7$ * * * *
Running the Bellman algorithm on Equation (6.2) produces a function that (barely) fails to be convex near the corners of $\Omega$, but the same is true for Method M2, see Figure 6. This example was also studied by Dean and Glowinski [10] (using a variational formulation) and Oberman [27] (wide stencil finite difference discretization), all with similar results. In [5], they suggest comparing the different methods using the minimum (i.e. the value at the origin) of the solution. The wide stencils methods from Oberman [27] are known to converge monotonically to a viscocity supersolution of Equation (2.1), and are thus too large. If the solution produced by the Bellman method had been convex, monotonicity would have implied that it too is too large, but since convexity fails near the corners of the square, we have no guarantee for this. Our method produces a solution which is slightly larger than the M2 and M1 methods, see Table 2.
@@ 6.3.2. Unbounded Monge–Ampère
If we instead do the computations on a slightly smaller square, e.g. $0,0.99]^{2}$, the three methods run at full accuracy ($\sim N^{-2}$), and the Bellman method runs 2–3 times faster than M2.
@@ 7. Conclusions
The Bellman algorithm relies heavily on the assumption that the solution to the Monge–Ampère equation is strictly convex on the domain. When this assumption is satisfied, the performance of the algorithm is very good. Using an $N\times N$ mesh, the discretization error in supremum and $\ell^{2}$ norm is proportional to $N^{-2}$, and is achieved with few iterations (typically fewer than $10$) resulting in a running time which is proportional to $N^{2.7}$$N^{2.8}$.